# YITP-25-92, IPMU25-0031

### v1

@@ ABSTRACT
We investigate observational constraints on cubic curvature corrections to general relativity by analyzing quasi-periodic oscillations (QPOs) in accreting black hole systems. In particular, we study Kerr black hole solution corrected by cubic curvature terms parameterized by $\beta_{5}$ and $\beta_{6}$. While $\beta_{6}$ corresponds to a field-redefinition invariant structure, the $\beta_{5}$ term can in principle be removed via a field redefinition. Nonetheless, since we work in the frame where the accreting matter minimally couples to the metric, $\beta_{5}$ is in general present. Utilizing the corrected metric, we compute the QPO frequencies within the relativistic precession framework. Using observational data from GRO J1655--40 and a Bayesian analysis, we constrain the coupling parameters to $|\beta_{5}|<(27.73\,\mathrm{km})^{4}$ and $|\beta_{6}|<(15.34\,\mathrm{km})^{4}$. These bounds improve upon existing constraints from big-bang nucleosynthesis and the speed of gravitational waves.
@@ 1 Introduction
Again the Ricci-flatness of the Kerr solution implies that only the last two terms in (4) provide nontrivial corrections to the Kerr metric [3, 4]. Within these two terms, only $\beta_{6}$ is invariant under field redefinitions, and thus it is the only term that contributes to black hole thermodynamics [3] and multipole moments, as we shall see below. We consider the most general parity even cubic curvature terms in the Lagrangian because we assume that the matter still minimally couples to the gravity. If one performs a field redefinition to remove $\beta_{1},\dots,\beta_{5}$ terms as considered in [5, 6, 7], the matter coupling to gravity will have to contain higher curvature terms. In particular, in the eikonal limit, a particle’s trajectory will deviate from the geodesics. We are aware that theoretical work concerning infrared causality has imposed bounds on the cutoff scale suppressing the higher curvature corrections [8, 9, 10], based on which whether EFT extensions of general relativity can have observational consequences has been questioned  [^2] The very recent work [11] suggested the opposite to the statement of [9].. In this work, we shall adopt a more pragmatic viewpoint similar to [7]. Causality-based arguments reflect the difficulty to construct viable EFT of quantum gravity. However, it is still valuable to provide independent constrains on the EFT coefficients purely based on observations, given the fact that even km-scale corrections are not ruled out by current experiments [7].
@@ 2 Kerr black hole solution with higher-derivative corrections
$+(1+H_{4})\Sigma\frac{dx^{2}}{1-x^{2}}+(1+H_{5})\Big{(}r^{2}+a^{2}+\frac{2\mu ra^{2}(1-x^{2})}{\Sigma}\Big{)}(1-x^{2})d\phi^{2}\ ,$
@@
where $H_{i}^{(n,p,k)}$ are constants and for each undetermined function, the number of $k_{\rm max}$ depends on $n$ and $p$. Substituting (7), (8) and (9) into equations of motion, we solve for $H_{i}^{(n,p,k)}$ order by order in $\chi$ up to ${\cal O}(\chi^{10})$. In the computation below, we shall utilize the solution up to ${\cal O}(\chi^{4})$ which is already rather lengthy. Thus we present the perturbative solution up to this order in an accompanying Mathematica notebook  [^3] Available at https://https://github.com/pangyi1/Kerr-Cubic.. In the data we will use, $\chi\sim 0.2-0.3$, which means our truncation of the perturbative solution at order ${\cal O}(\chi^{4})$ should be good enough for the precision of the data. We have chosen the parameters $\mu$ and $\chi$ properly so that up to linear order in $\beta_{i}$, the mass and angular momentum remain uncorrected
@@
We utilize the data available for GRO J1655--40 to constrain $\beta_{5}$ and $\beta_{6}$. The detection of QPOs from this source was achieved through observations with RXTE [33, 34]. Using X-ray timing techniques, three distinct QPO frequencies have been identified [35, 22]. There are also other X-ray binaries like XTE J1550--564. However, for this system two distinctive peaks are observed [36, 37]. Our framework applies to X-ray binaries with three simultaneous peaks. The relevant data are summarized in Table 1. To perform our Bayesian analysis, we define the $\chi$-square from QPOs as
@@
Table 1: Parameters $\nu_{U}$, $\nu_{L}$ and $\nu_{C}$ for GRO J1655--40, where $\nu_{C}$, $\nu_{L}$ and $\nu_{U}$ are observed values for $\nu_{n}$, $\nu_{p}$ and $\nu_{\phi}$ respectively.
Additionally, we consider non-informative uniform priors as $10<a/M<$ 1, $104<r/M<$ 10, $-50<\beta_{5}/(5M_{\odot})^{4}<50$ and $-50<\beta_{6}/(5M_{\odot})^{4}<50$. The mass for this system has been determined from ellipsoidal light-curve fits to the quiescent B, V , R and I light curves [38]. We use this uncertainty assuming a Gaussian prior for the mass as
@@
Figure 1: Two dimensional and one dimensional marginal distributions for $\beta_{5}$. The contours show 68%, 95% and 99% credible intervals. Dashed lines are the 5%, 50%, and 95% percentiles of the distribution. At one-$\sigma$ we have $-12.15<\frac{\beta_{5}}{(5M_{\odot})^{4}}<9.31$.
Figure 2: Two dimensional and one dimensional marginal distributions for $\beta_{6}$. The contours show 68%, 95% and 99% credible intervals. Dashed lines are the 5%, 50%, and 95% percentiles of the distribution. At one-$\sigma$ we have $-1.17<\frac{\beta_{6}}{(5M_{\odot})^{4}}<0.25$.
The results are illustrated in figures 1 and 2 using Dynesty [39]. The plots show noticeable correlations between $\beta_{5}$ and other parameters such as $\chi$ and $M$. We also find that $\beta_{6}$, $\chi$ and $M$ are highly correlated. This hinders our ability to constrain them with high precision. For $\beta_{6}$ we manage to find tighter bounds compared to $\beta_{5}$. We find that $-1.17<\frac{\beta_{6}}{(5M_{\odot})^{4}}<0.25$ at one-$\sigma$ level. This corresponds to $|\beta_{6}|<(15.34\,{\rm km})^{4}$. For $\beta_{5}$ we have $-12.15<\frac{\beta_{5}}{(5M_{\odot})^{4}}<9.31$, translating to $|\beta_{5}|<(27.73\;{\rm km})^{4}$. For reference, Schwarzschild radius for this source is $r_{s}\approx 32.45\;{\rm km}$.
@@ 4 Discussion and Conclusion
Our bound from QPOs is stronger than these bounds from the big-bang nucleosynthesis and the speed of gravitational waves. Our bound is also consistent with the causality bound [8]  [^4] The causality bound used here is derived from the requirement that the correction to the time delay caused by the higher curvature term is negative and unsolvable, namely $-\delta T>1/\omega$, where $\delta T$ is the correction to time delay, and $\omega$ is the frequence of the signal. In the recent work [11], it is argued that the causality bound is not better than experimental constraints. There is no contradition between [8] and [11], because the latter considered the modification to GR is exact rather than being perturbative correction and required the total time delay be negative. ,
@@ Acknowledgements
This work of LM and YP is supported by the National Key Research and Development Program No. 2022YFE0134300. L.M. is also supported in part by National Natural Science Foundation of China (NSFC) grant No. 12447138, Postdoctoral Fellowship Program of CPSF Grant No. GZC20241211 and the China Postdoctoral Science Foundation under Grant No. 2024M762338. The work of SM was supported in part by JSPS (Japan Society for the Promotion of Science) KAKENHI Grant No. JP24K07017 and World Premier International Research Center Initiative (WPI), MEXT, Japan.
Appendix A Geometric identities
@@
00 0 $=$ $-12RR_{\mu\nu}R^{\mu\nu}+R^{3}+24R_{\mu\lambda\nu\rho}R^{\mu\nu}R^{\lambda\rho}+16R_{\mu}^{\nu}R_{\nu}^{\lambda}R_{\lambda}^{\mu}-24R_{\mu\nu}R^{\mu\lambda\rho\sigma}R^{\nu}_{\ \lambda\rho\sigma}$
@@
00 0 $=$ $4R_{\mu}^{\nu}R_{\nu}^{\lambda}R_{\lambda}^{\mu}-4RR_{\mu\nu}R^{\mu\nu}+{\textstyle{\frac{\scriptstyle 1}{\scriptstyle 2}}}R^{3}+4R_{\mu\lambda\nu\rho}R^{\mu\nu}R^{\lambda\rho}+{\textstyle{\frac{\scriptstyle 1}{\scriptstyle 2}}}RR^{\mu\lambda\rho\sigma}R_{\mu\lambda\rho\sigma}$ (30)

### v2

@@ ABSTRACT
We investigate observational constraints on cubic curvature corrections to general relativity by analyzing quasi-periodic oscillations (QPOs) in accreting black hole systems. In particular, we study Kerr black hole solution corrected by cubic curvature terms parameterized by $\beta_{5}$ and $\beta_{6}$. While $\beta_{6}$ corresponds to a field-redefinition invariant structure, the $\beta_{5}$ term can in principle be removed via a field redefinition. Nonetheless, since we work in the frame where the accreting matter minimally couples to the metric, $\beta_{5}$ is in general present. Utilizing the corrected metric, we compute the QPO frequencies within the relativistic precession framework. Using observational data from GRO J1655-40 and a Bayesian analysis, we constrain the coupling parameters to $12.31<\frac{\beta_{5}}{(5M_{\odot})^{4}}<24.15$ and $1.99<\frac{\beta_{6}}{(5M_{\odot})^{4}}<0.30$ at 2-$\sigma$. These bounds improve upon existing constraints from big-bang nucleosynthesis and the speed of gravitational waves.
@@ 1 Introduction
Again the Ricci-flatness of the Kerr solution implies that only the last two terms in (4) provide nontrivial corrections to the Kerr metric [3, 4]. Within these two terms, only $\beta_{6}$ is invariant under field redefinitions, and thus it is the only term that contributes to black hole thermodynamics [3] and multipole moments, as we shall see below  [^2] In [5] an action-based order-reduction procedure was applied to the $\beta_{6}$ term to explicitly remove the spurious degrees from the EFT. The resulting effective action and equations of motion are expected to be useful e.g. for establishing a well-posed formulation of the numerical evolution of the system. . We consider the most general parity even cubic curvature terms in the Lagrangian because we assume that the matter still minimally couples to the gravity. If one performs a field redefinition to remove $\beta_{1},\dots,\beta_{5}$ terms as considered in [6, 7, 8], the matter coupling to gravity will have to contain higher curvature terms. In particular, in the eikonal limit, a particle’s trajectory will deviate from the geodesics. We are aware that theoretical work concerning infrared causality has imposed bounds on the cutoff scale suppressing the higher curvature corrections [9, 10, 11], based on which whether EFT extensions of general relativity can have observational consequences has been questioned  [^3] The very recent work [12] suggested the opposite to the statement of [10].. In this work, we shall adopt a more pragmatic viewpoint similar to [8]. Causality-based arguments reflect the difficulty to construct viable EFT of quantum gravity. However, it is still valuable to provide independent constraints on the EFT coefficients purely based on observations, given the fact that even km-scale corrections are not ruled out by current experiments [8].
@@ 2 Kerr black hole solution with higher-derivative corrections
$+(1+H_{5})\Sigma\frac{dx^{2}}{1-x^{2}}+(1+H_{4})\Big(r^{2}+a^{2}+\frac{2\mu ra^{2}(1-x^{2})}{\Sigma}\Big)(1-x^{2})d\phi^{2}\ ,$
@@
where $H_{i}^{(n,p,k)}$ are constants and for each undetermined function, the number of $k_{\rm max}$ depends on $n$ and $p$. Substituting (7), (8) and (9) into equations of motion, we solve for $H_{i}^{(n,p,k)}$ order by order in $\chi$ up to ${\cal O}(\chi^{10})$. In the computation below, we shall utilize the solution up to ${\cal O}(\chi^{4})$, which is already quite lengthy. Thus we present the perturbative solution up to this order in an accompanying Mathematica notebook  [^4] Available at https://https://github.com/pangyi1/Kerr-Cubic. In the solution, we have chosen parameters $\mu$ and $\chi$ such that the mass and angular momentum do not depend on $\beta_{5}$ and $\beta_{6}$ at linear order. By contrast, the solution given in [7] did not make such a choice. This is the main reason why our solution looks different from the solution there. . In the data we will use, $\chi\sim 0.2-0.3$, which means our truncation of the perturbative solution at order ${\cal O}(\chi^{4})$ should be good enough for the precision of the data. We have chosen the parameters $\mu$ and $\chi$ properly so that up to linear order in $\beta_{i}$, the mass and angular momentum remain uncorrected
@@
We utilize the data available for GRO J1655-40 to constrain $\beta_{5}$ and $\beta_{6}$. The detection of QPOs from this source was achieved through observations with RXTE [34, 35]. Using X-ray timing techniques, three distinct QPO frequencies have been identified [36, 23]. There are also other X-ray binaries like XTE J1550-564. However, for this system two distinctive peaks are observed [37, 38]. Our framework applies to X-ray binaries with three simultaneous peaks. The relevant data are summarized in table.1.
To perform our Bayesian analysis, we define the $\chi$ square from QPOs as
@@
Table 1: Parameters $\nu_{U}$, $\nu_{L}$ and $\nu_{C}$ for GRO J1655-40, where $\nu_{C}$, $\nu_{L}$ and $\nu_{U}$ are observed values for $\nu_{n}$, $\nu_{p}$ and $\nu_{\phi}$ respectively. The errors show 68% confidence intervals obtained after fitting Lorentzian shapes to the power density spectrum using the XSPEC package [36].
Additionally, we consider non-informative uniform priors as $0<a/M<1$, $4<r/M<10$, $50<\beta_{5}/(5M_{\odot})^{4}<50$ and $50<\beta_{6}/(5M_{\odot})^{4}<50$. The mass for this system has been determined from ellipsoidal light-curve fits to the quiescent B, V , R and I light curves [39]. We use this uncertainty assuming a Gaussian prior for the mass as
@@
Figure 1: Two dimensional and one dimensional marginal distributions for $\beta_{5}$. The contours show 68%, 95% and 99% credible intervals. Dashed lines are the 5%, 50%, and 95% percentiles of the distribution. At 2-$\sigma$ we have $12.31<\frac{\beta_{5}}{(5M_{\odot})^{4}}<24.15$.
Figure 2: Two dimensional and one dimensional marginal distributions for $\beta_{6}$. The contours show 68%, 95% and 99% credible intervals. Dashed lines are the 5%, 50%, and 95% percentiles of the distribution. At 2-$\sigma$ we have $1.99<\frac{\beta_{6}}{(5M_{\odot})^{4}}<0.30$.
The results are illustrated in figures 1 and 2 using Dynesty [40]. To run computations in parallel, we used the pool function in Dynesty to speed up the computations. The plots show noticeable correlations between $\beta_{5}$ and other parameters such as $\chi$ and $M$. We also find that $\beta_{6}$, $\chi$ and $M$ are highly correlated. This hinders our ability to constrain them with high precision. For $\beta_{6}$ we manage to find tighter bounds compared to $\beta_{5}$. We find that $1.99<\frac{\beta_{6}}{(5M_{\odot})^{4}}<0.30$ at 2-$\sigma$ level. For $\beta_{5}$ we have $12.31<\frac{\beta_{5}}{(5M_{\odot})^{4}}<24.15$ at 2-$\sigma$ level.
@@ 4 Discussion and Conclusion
Our bound from QPOs is stronger than these bounds from the big-bang nucleosynthesis and the speed of gravitational waves. Our bound is also consistent with the causality bound [9]  [^5] The causality bound used here is derived from the requirement that the correction to the time delay caused by the higher curvature term is negative and unsolvable, namely $\delta T>1/\omega$, where $\delta T$ is the correction to time delay, and $\omega$ is the frequence of the signal. In the recent work [12], it is argued that the causality bound is not better than experimental constraints. There is no contradition between [9] and [12], because the latter considered the modification to GR is exact rather than being perturbative correction and required the total time delay be negative. ,
@@ Acknowledgements
This work of LM and YP is supported by the National Key Research and Development Program No. 2022YFE0134300, the National Natural Science Foundation of China (NSFC) under Grant No. 12175164 and Tianjin University Self-Innovation Fund Extreme Basic Research Project Grant No. 2025XJ21-0007. L.M. is also supported in part by National Natural Science Foundation of China (NSFC) grant No. 12447138, Postdoctoral Fellowship Program of CPSF Grant No. GZC20241211 and the China Postdoctoral Science Foundation under Grant No. 2024M762338. The work of SM was supported in part by JSPS (Japan Society for the Promotion of Science) KAKENHI Grant No. JP24K07017 and World Premier International Research Center Initiative (WPI), MEXT, Japan.
Appendix A Geometric identities and equations of motion
@@
$0$ $=$ $-12RR_{\mu\nu}R^{\mu\nu}+R^{3}+24R_{\mu\lambda\nu\rho}R^{\mu\nu}R^{\lambda\rho}+16R_{\mu}^{\nu}R_{\nu}^{\lambda}R_{\lambda}^{\mu}-24R_{\mu\nu}R^{\mu\lambda\rho\sigma}R^{\nu}_{\ \lambda\rho\sigma}$
@@
$0$ $=$ $4R_{\mu}^{\nu}R_{\nu}^{\lambda}R_{\lambda}^{\mu}-4RR_{\mu\nu}R^{\mu\nu}+{\textstyle{\frac{\scriptstyle 1}{\scriptstyle 2}}}R^{3}+4R_{\mu\lambda\nu\rho}R^{\mu\nu}R^{\lambda\rho}+{\textstyle{\frac{\scriptstyle 1}{\scriptstyle 2}}}RR^{\mu\lambda\rho\sigma}R_{\mu\lambda\rho\sigma}$ (30)
@@
The equation of motion used in our analysis is given as
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ $=$ $-P_{(\mu}^{\ \ \alpha\beta\gamma}R_{\nu)\alpha\beta\gamma}-2\nabla^{\alpha}\nabla^{\beta}P_{(\mu|\alpha|\nu)\beta}$ (31)
$+$ $\frac{1}{2}g_{\mu\nu}\left(\beta_{5}RR^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}+\beta_{6}R^{\mu\nu}_{\ \ \lambda\rho}R^{\lambda\rho}_{\ \ \sigma\delta}R^{\sigma\delta}_{\ \ \mu\nu}\right)\,$,
where we define the tensor $P_{\mu\nu\rho\sigma}$
$P_{\mu\nu\rho\sigma}$ $=$ $\beta_{5}\Big[\frac{1}{2}(g_{\mu\rho}g_{\nu\sigma}-g_{\mu\sigma}g_{\nu\rho})R_{\alpha\beta\gamma\lambda}R^{\alpha\beta\gamma\lambda}+2RR_{\mu\nu\rho\sigma}\Big]+3\beta_{6}R_{\mu\nu}^{\ \ \alpha\beta}R_{\rho\sigma\alpha\beta}\,.$ (32)
The right hand side of (31) evaluated on the uncorrected Kerr black hole serves as the effective stress tensor for the the perturbations at first order in $\beta_{i}$. The other terms in the effective Lagrangian (4) are irrelevant in our discussion precisely because their corresponding effective stress tensor vanish on the uncorrected Kerr black hole solution.